OpenStudy (anonymous):

Graph the relation {(5, 0), (0, 5), (5, 1), (1, 5)}. Is it a function? Why or why not?

OpenStudy (anonymous):

if it's a function, that means that every input hads only one output. ususally, the first number in the ordered pair is the input and the second is the output.

OpenStudy (anonymous):

y is not a function of x because two different x's give different y's ((5,0) and (5,1) if x is 5 then y can't both be 0 and 1. So I'd say no not a function.

OpenStudy (anonymous):

these are my choices. A.Yes, it passes the horizontal-line test. B. No, it does not pass the horizontal-line test C. Yes, it passes the vertical-line test. D.No, it does not pass the vertical-line test..

OpenStudy (anonymous):

@Drahgnus please read the code of conduct. don't just give out answers. instead help lead them to a place of understanding.

OpenStudy (anonymous):

doesn't pass the vertical line test because at x = 1 there are two y's above it. Btw pgpilot archer vice is awesome

OpenStudy (anonymous):

oh lol didn't know that sorry first 20 minutes on this site

OpenStudy (anonymous):

yeah, i can't wait!

OpenStudy (anonymous):

lol its ok. so my answer would be C.??

OpenStudy (anonymous):

no worries, how did you get a score of 42 in 10 minutes?

OpenStudy (anonymous):

nope answer is not C.

OpenStudy (anonymous):

ohh...? does it pass the vertical??

OpenStudy (anonymous):

so snap so it would be D. my bad

OpenStudy (anonymous):

you tell me, what does it mean to pass the vertical line test?

OpenStudy (anonymous):

I've answered like 20 questions they're all kinda simple here, I'm trying to get an answer to differential equation question. Look that type of math up and thats how :P Aerospace engineering ;) Imagine putting a vertical line at x = 1, does it hit more than one y when you do that?

OpenStudy (anonymous):

funny, i have a degree in math and didn't get points that quick.

OpenStudy (anonymous):

what kind of diffy q?

OpenStudy (anonymous):

I wish I could both of you guys medals.. sorry I cant though

OpenStudy (anonymous):

Give it to pg

OpenStudy (anonymous):

did you get your question figured out, @PeytonPerez372 ?

OpenStudy (anonymous):

no, i don't care. give it to Drahgnus

OpenStudy (anonymous):

sorry guys you both were a big help though

OpenStudy (anonymous):

that's why we're here

OpenStudy (anonymous):

RAMPAAAAAGEEEEEE

OpenStudy (anonymous):

LANA!!!!

OpenStudy (anonymous):

SPLOOOSH

OpenStudy (anonymous):

Bloody mary, blessed are you among wingspantails. Pray for me now in the hour of my death, which I hope is soon. Amen.

OpenStudy (anonymous):

oh for my differential equation.. you have a degree in math? that might help... okay so its a homogenous differential, so youre supposed to use u = y/x y' = y/x + x / [(x^2 * y)^1/3 + (xy)^1/2 + x ]

OpenStudy (anonymous):

i've literally waited my whole life to say these words... I'm comandeering this AIRBOAT!!!

OpenStudy (anonymous):

hahahaha who do i remind you of?! burt reynolds in what?!

OpenStudy (anonymous):

okay and what did you get after subbing?

OpenStudy (anonymous):

GATOR

OpenStudy (anonymous):

oh wait, the one before had the airboats... WHITE LIGHTNING!

OpenStudy (anonymous):

I can't get it in a form of y/x that's my problem, I know in the end the y' = u + xu' and then the right side the first part is u + ... its just a mess I don't even know, I've tried dividing by xy^1/2 but it didn't look promising.. tried dividing by y.. just not working for me :\

OpenStudy (anonymous):

okay, hold on a sec...

OpenStudy (anonymous):

and I think the reason I received points so fast was because I just gave them the answer, didn't show any work and did that to like 30 simple problems lol.

OpenStudy (anonymous):

thanks a lot :D if you want I can link you to my question after so you can get a medal

OpenStudy (anonymous):

so did you get \[y'=\frac{u'x^2+y}{x}\]

OpenStudy (anonymous):

you guys are like trolling my question lol

OpenStudy (anonymous):

sorry this should be a link to my question http://openstudy.com/users/drahgnus#/updates/52d979fae4b05ee05437f785

OpenStudy (anonymous):

sorry @PeytonPerez372